OpenStudy (anonymous):

Please help!!! What is the lateral area of a rectangular prism if the base edges are 8 feet and 6 feet and the height is 10 feet?

OpenStudy (anonymous):

I have no idea how to do this...

jimthompson5910 (jim_thompson5910):

use the formula LSA = 2*(L*H + W*H) where LSA = lateral surface area L = length W = width H = height

jimthompson5910 (jim_thompson5910):

in this case LSA = unknown (we're solving for this) L = 8 W = 6 H = 10

OpenStudy (anonymous):

so 2*(L*H + W*H), 2*(8*10+6*10), 2*(80+60) 2*140, 280 ?

jimthompson5910 (jim_thompson5910):

yep LSA = 2*(L*H + W*H) LSA = 2*(6*10 + 8*10) LSA = 2*(60 + 80) LSA = 2*(140) LSA = 280

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome